OpenStudy (anonymous):

given the equation y-4 =3/4(x+8) in point -slope form , identify the equation of the same line in standard form.

OpenStudy (anonymous):

do you know the standard line equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y = mx +c, where m is your gradient, c is your intercept and x and y are x and y

OpenStudy (anonymous):

so can you take your equation and rearrange it to make it y = something?

OpenStudy (anonymous):

can u do that for me cuz i dont know how

OpenStudy (anonymous):

I will try and help you understand so that you can do it, So you start with this right:\[y -4 = \frac{ 3 }{ 4 }(8+x)\] Can you expand the brackets on the right hand side?

OpenStudy (anonymous):

sorry but kinda in a hurry can u give answer plz

OpenStudy (***[isuru]***):

okiii from the point slope form... y-4 =3/4(x+8) we can directly get the slope... here the slope is... 3/4 also.. we can get the co-ordinate of a point in the line which is.. (-8 , 4) now the slandered notation is y = mx + c hence we know the slope we can write it as.. y = 3/4x + c now.. let's substitute the co-ordinates to the function.. 4 = (3/4)(-8) +c c = 10 so..the slandered form is y = (3/4)x + 10

OpenStudy (anonymous):

oh thx

OpenStudy (***[isuru]***):

u r welcome!!